OpenStudy (anonymous):

Three golfers, Vijay, Ernie, and Phil, during 5 years had a total of 196 top-10 finishes. Ernie had 11 more top-10 finishes than Phil, Vijay had 14 more top-10 finishes than Phil. Determine the number of top-10 finishes for each golfer. Can someone explain this to me? Step-by-Step through the matrix. and how to set it up on paper to enter it into the matrix.

OpenStudy (anonymous):

v + E + P = 196 E = P + 11 V = P + 14 Substitute and solve.

OpenStudy (anonymous):

That doesnt help me one bit, i already had that. I dont know what to do after.

OpenStudy (anonymous):

You mean you don't know how to solve?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

V + E + P = 196 (P + 14) + (P + 11) + P = 196 3P + 25 = 196 3P = 171 P = 171/3 = 57 P = 57 E = 57 + 11 V = 57 + 14

OpenStudy (anonymous):

I accidentally been putting a 10 instead of 11. cause the way it was worded. lol. Thank you!